OpenStudy (anonymous):

Find the mean, median, mode, and range of the data set. Daily temperature in degrees Celsius: 24, 24, 25, 27, 31, 32, 38, 39 Options: mean = 24, median = 30, mode = 29, range = 16 mean = 29, median = 30, mode = 24, range = 15 mean = 30, median = 29, mode = 24, range = 15 mean = 30, median = 24, mode = 29, range = 16

OpenStudy (anonymous):

3rd one down

OpenStudy (anonymous):

Agreed its C.

OpenStudy (anonymous):

I realize your question has been answered, however, they are just answers...let me show you how to do these problems... The mean is the average... first you have to add the numbers... 24 + 24 + 25 + 27 + 31 + 32 + 38 + 39 = 240 Now take your total and divide it by how many numbers there are.. 240/8 = 30 (this is your mean) The median is the middle number... Since there is 8 numbers, you will have to take the two middle numbers and add them, then divide by 2 (just like finding the average) So, the two middle numbers are 27 and 31. added = 58. Now divide it by 2 and you get 29 The mode is the number that appears the most. Just letting you know, there doesn't always have to be a mode, but there can also be more then one mode. In this problem, the mode is 24 because it appears twice in the set of numbers and no other number does. The range is the difference between the highest and the lowest number... So, the range is 39 - 24 = 15 The answer is C....you probably already knew that because of the previous answers, but now, you have a little bit of understanding as how to do it.